### v4

# Quotient Space Quantum Codes

Abstract
Additive codes and some nonadditive codes use the single and multiple invariant subspaces of the stabilizer G to construct quantum codes, respectively, so the selection of invariant subspaces is a key issue. In this letter, I provide the necessary and sufficient conditions for this problem and, for the first time, establish the quotient space codes to construct quantum codes. This new code unifies additive codes and codeword stabilized codes and can transmit classical codewords. New bounds for quantum codes are presented also, and a simple proof of the quantum Singleton bound is provided. The quotient space approach offers a concise and clear mathematical form for the study of quantum error-correcting codes.
preprint: APS/123-QED
Quantum error-correcting codes [1] are a coding scheme used to protect quantum information from noise and errors, and are a crucial problem for the realization of quantum computing and quantum communication [2, 3]. Since Shor [4] constructed the first code, a large number of coding schemes have been proposed, among which stabilizer codes or additive codes [5, 6, 7] are the most representative class of codes. They use a $2^{k}$-dimensional invariant subspace of an additive subgroup $G$ (stabilizer) of the quantum error operator group $E$ to construct codes. Additive codes are usually denoted as $[[n,k,d]],$ indicating that $n$ physical qubits are encoded as $k$ logical qubits and can correct $\lfloor\frac{d-1}{2}\rfloor$-qubit errors, where $d$ is the minimum distance. Another type of code is nonadditive codes, first proposed by Rains et al.[8], some of which use $L$ one-dimensional invariant subspaces of the stabilizer to construct codes, denoted as $((n,L,d)),$ representing $L$-dimensional subspaces in an $n$-dimensional Hilbert space, that can correct $\lfloor\frac{d-1}{2}\rfloor$qubit errors. The references [9, 10, 11] used different methods to construct nonadditive codes that are better than additive codes. Reference [12] created codeword stabilized codes (CWS) that unify additive and some good nonadditive codes. These methods are elegant but seem too esoteric and lack a language or concepts to succinctly express the problem, such as how to construct codes for $d\geq 4.$ This letter utilizes normed quotient spaces to establish quotient space codes. If a quotient space code $(n-k,L,d)$ that satisfies the measurement conditions exists, then there are $L$ additive codes of dimension $2^{k}$ that together form a quotient space quantum code $((n,2^{k}L,d)).$ When $L=1,$ an additive code is obtained, and when $k=0,$ a CWS code is obtained. The circuit design and decoding scheme of this code can be improved based on the additive code. It is observed that although CWS codes can include additive codes, these two code families have fundamentally different error correction methods. I also provide new bounds—measurement bounds, which reflect the unique properties of quantum codes compared to classical codes and distinguish between degenerate and nondegenerate codes. The choice of coset representatives and norms greatly simplifies the problem description, as shown in the proof of the quantum Singleton bound. I present four methods for constructing the code $((8,8,3))$ and provide an example of a degenerate CWS code $((12,2,5)).$ Preparation  First, I introduce the basics and some notations of quantum codes. The quantum error group $E_{n}$ is written as $E_{n}=\{i^{q}X(a)Z(b)\mid 0\leq q\leq 3,~{}a,~{}b\in\mathbb{F}_{2}^{n}\},$ and in situations where there is no confusion it is also written as $E.$ We define the mapping $\varphi:E_{n}\rightarrow\mathbb{F}_{2}^{2n},~{}e=i^{q}X(a)Z(b)\longmapsto\varphi(e)=(a|b),$ which induces a group isomorphism $\bar{\varphi}:\overline{E}_{n}=E_{n}/\{\pm 1,\pm i\}\rightarrow V=\mathbb{F}_{2}^{2n},~{}\bar{\varphi}(\bar{e})=(a|b),$ thus allowing quantum errors to be represented by vectors in $\mathbb{F}_{2}^{2n},$ and we may denote $\bar{e}=\varphi(e).$ Let $w_{Q}(\bar{e})$ be the quantum weight of $e,$ $w_{H}(\bar{e})$ be the Hamming weight of $e,$ $E(d)$ be the quantum errors with weight not exceeding $d,$ and $\overline{E}(d)$ be its image. The symplectic inner product of vectors in $\mathbb{F}_{2}^{2n}$ is defined as $(c_{1},c_{2})_{s}=a\textperiodcentered b^{\prime}+a^{\prime}\textperiodcentered b,~{}c_{1}=(a|b),~{}c_{2}=(a^{\prime}|b^{\prime})\in\mathbb{F}^{2n}_{2}.~{}C$ is a subspace of $\mathbb{F}^{2n}_{2},$ and $C^{\bot}_{s}=\{c_{s}\mid(c_{s},c)_{s}=0,~{}\forall~{}c\in C\},$ is the symplectic orthogonal space of $C.$ $G$ is an additive subgroup of $E$ and $\overline{G}=C,~{}G^{\bot}_{s}=\{e\mid e\in E_{n},~{}\bar{e}\in C^{\bot}_{s}\},d_{m}=min\{w_{Q}(c)\mid c\in(C^{\bot}_{s})\backslash C\},~{}C(d)=\overline{E}(d)\bigcap C,~{}C(d)^{\bot}_{s}$ is the symplectic orthogonal space of $C(d),$ and it is known that $C^{\bot}_{s}\subset C(d)^{\bot}_{s}.$ A characteristic of a finite commutative group $G$ is a group homomorphism from $G$ to the nonzero complex multiplication group $\mathbb{C}^{*}$
$\lambda:G\rightarrow\mathbb{C}^{*},~{}\lambda(gg^{\prime})=\lambda(g)\lambda(g^{\prime})~{}(g,~{}g^{\prime}\in G).$
The set of all characteristics of $G$ forms a group $\widehat{G},$ and the characteristic group of $C$ is denoted as $\widehat{C}.$ Characterization of Invariant Subspaces   To express the problem in concise mathematical form, it is necessary to characterize the invariant subspaces of $G$ and the action of quantum errors on them. Due to the isomorphism between a finite abelian group $G$ and its character group $\widehat{G}$ [13], and the natural isomorphism between the additive character group $\widehat{C}$ of $C$ and the additive quotient group $F_{2}^{2n}/C^{\bot}_{s}$ [14], it is easy to prove using the homomorphism theorem that
$\widehat{G}\cong G\cong C\cong\widehat{C}\cong\dfrac{\mathbb{F}^{2n}_{2}}{C^{\bot}_{s}}\cong\dfrac{E_{n}}{G^{\bot}_{s}}.$ (1)
Let $I=\frac{\mathbb{F}^{2n}_{2}}{C^{\bot}_{s}},$ and $[i]$ be a coset representative for $i,$ therefore, we have
$\widehat{G}=\widehat{C}=\{\lambda_{[i]}\mid[i]\in I\}.$ (2)
Although the notation for these two types of character groups is the same, we can distinguish them based on the objects they act on. In particular, $\lambda_{[i]}(g)=\lambda_{i}(g)=(-1)^{(i,\bar{g})_{s}}~{}(g\in G,~{}i\in\mathbb{F}^{2n}_{2}),~{}\lambda_{[i]}(c)=\lambda_{i}(c)=(-1)^{(i,c)_{s}}~{}(c\in C),$ and we also denote $\lambda_{e}(g)=\lambda_{\bar{e}}(\bar{g})$ as the character of $G.$ It can be observed that $\lambda_{i}=\lambda_{j}\Leftrightarrow[i]=[j],$ and $\lambda_{e_{1}}=\lambda_{e_{2}}\Leftrightarrow[\overline{e_{1}}]=[{\overline{e_{2}}}].$ Since an $n$-dimensional Hilbert space has an orthogonal direct sum decomposition [14], $\mathbb{C}^{2^{n}}=\mathop{\oplus}_{\lambda\in\widehat{G}}Q(\lambda),$ where each $Q(\lambda)=\{v\in\mathbb{C}^{2^{n}}\mid~{}\forall~{}g\in G,~{}g(v)=\lambda(g)v\}$ is an invariant subspace of $G$ and can be used to construct additive codes. Let $Q(\lambda_{[i]})=Q([i])=Q(i),$ then $Q(i)=Q(j)\Leftrightarrow[i]=[j],$ therefore, the above decomposition can be written as
$\mathbb{C}^{2^{n}}=\mathop{\oplus}_{[i]\in I}Q([i]).$ (3)
Let $Q(I)=\{Q([i])\mid[i]\in I\},$ the action of quantum errors on the invariant subspaces was described by
Theorem 1
The action of $E$ on $Q(I)$ corresponds to the additive of I on itself, and $G^{\bot}_{s}$ corresponds to $[\mathbf{0}].$
Proof  For any $|v\rangle\in Q([i]),$ and $g\in G,$ we have
$g(e|v\rangle)$ $=(ge)|v\rangle=(-1)^{(\bar{e},\bar{g})_{s}}(eg)|v\rangle=(-1)^{(\bar{e},\bar{g})_{s}}e(g|v\rangle)$
$=\lambda_{e}(g)e\lambda_{[i]}(g)|v\rangle=\lambda_{[\overline{e}]}(g)\lambda_{[i]}(g)e|v\rangle$
$=\lambda_{[\overline{e}]+[i]}(g)e|v\rangle,$
where $\lambda_{[\overline{e}]+[i]}$ denotes the character associated with the coset $[\overline{e}]+[i].$ Therefore, $e|v\rangle\in Q([\overline{e}]+[i]),$ which means that the action $eQ([i])=Q([j])$ corresponds to $[\overline{e}]+[i]=[j],$ and vice versa. $\square$ Quotient Space Codes   So, the problem of selecting invariant subspaces was transformed into the problem of selecting cosets, which naturally leads to the concept of the quotient space codes. To establish a quotient space code, we first define the distance and norm of the quotient space. For any $[x],~{}[y]\in I,$ we define the distance as
$d({[x],[y]})=\inf_{\tilde{x}\in[x],\tilde{y}\in[y]}d(\tilde{x},\tilde{y}),$ (4)
and define the norm as
$\|[x]\|=\mathop{\inf}\limits_{\tilde{x}\in[x]}\|\tilde{x}\|.$ (5)
When the distance satisfies the translation invariance property, i.e., $d(x+z,y+z)=d(x,y),$ the above definitions satisfy the axioms of a metric and a norm, respectively [15]. To ensure that $\|x\|=w_{Q}(x)$ becomes a norm, I define $|\alpha|=1$ if $\alpha\neq 0$ ($\alpha\in\mathbb{F}_{q}$). This allows the quantum weight to satisfy the homogeneity axiom $\|\alpha x\|=|\alpha|\|x\|,$ which is consistent with the definition of a norm. I call this the quotient minimum norm. Similarly, the same treatment can be applied to the Hamming weight. Using $d(x,y)=\|x-y\|,$ the distance and norm can be mutually converted, and thus $I$ becomes a metric space and a normed space. To extend the conclusions of this letter to quantum codes over $q$-ary fields[16, 17], here is a general definition of a quotient space code. Let $V=\mathbb{F}_{q}^{2n},$ $H$ be a subspace of $V$ and $W=V/H$ be the normed (or metric) quotient space of $V$ with respect to $H.$ If $dimW=n-k,$ then any non-empty subset $\Omega$ of $W$ is called a quotient space code ($\mathbf{QSC}$) (to avoid confusion with quantum spherical codes, it can also be called a classical QSC or coset code). Each element (coset) in $\Omega$ is called a codeword. If $|\Omega|=L,$ and $d$ is the minimum distance between cosets in $\Omega,$ then $\Omega$ is called an $(n-k,L,d)_{q}$ QSC, where $n-k$ represents the code length and $L$ represents the number of codewords. Let $l=\log_{q}L$ be the number of information bits in the code $\Omega,$ $\Omega$ can also be denoted as $[n-k,l,d]_{q}.$ Similar to classical error-correcting codes [18], a QSC with a distance of $d$ can detect quotient errors up to $d-1,$ and can correct quotient errors up to $\lfloor\frac{d-1}{2}\rfloor.$ New Codes  The preliminary work is complete, and I now present the new code.
Theorem 2
Let $C$ be a symplectic self-orthogonal subspace in $\mathbb{F}_{2}^{2n},$ where $dimC=n-k$ with $0\leq k\leq n,$ let $I=\mathbb{F}_{2}^{2n}/{C_{s}^{\bot}},$ and $\Omega$ be a QSC in $I$ with parameters $(n-k,L,d),$ where $d\leq d_{m}$ and $\Omega$ belongs to the same coset in the quotient space $\mathbb{F}_{2}^{2n}/C(d-1)_{s}^{\bot}.$Then, there exists a quantum code $Q(\Omega)((n,2^{k}L,d)),$ where
$Q(\Omega)=\bigoplus_{[i]\in\Omega}Q([i]).$ (6)
Proof Because $C\subseteq C_{s}^{\bot},$ we can lift $C$ to $E_{n}$ and obtain a stabilizer $G$ such that $\overline{G}=C.$ First, let us prove that $dim(Q(\Omega))=2^{k}L.$ From equation (3), we know that $\mathbb{C}^{2^{n}}=\bigoplus_{[i]\in I}Q([i]).$ Since the action of $I$ on $I$ is transitive, according to Theorem 1, each $Q([i])$ is isomorphic and has the same dimension. Hence, $dim(Q([i]))=\frac{2^{n}}{|I|}=\frac{2^{n}}{2^{n-k}}=2^{k},$ implying $dim(Q(\Omega))=dim(Q([i]))|\Omega|=2^{k}L.$ Next, I prove that the minimum distance of $Q(\Omega)$ is $d.$ Let $|v_{i}\rangle,|v_{j}\rangle\in Q(\Omega),$ where $|v_{i}\rangle\in Q([i])$ and $|v_{j}\rangle\in Q([j]),$ with $\langle v_{i}|v_{j}\rangle=0.~{}$Knill and Laflamme [19] and Bennett et al [20] have shown that a quantum code can correct errors $e\in E_{n}(d-1)$ iff
$\langle v_{i}|e|v_{j}\rangle=f(e)\delta_{ij},$ (7)
where $f(e)$ depends only on $e.$ I prove this condition separately for the cases where $\delta_{ij}=0$ and $\delta_{ij}=1.$ When $\delta_{ij}=0,$ the necessary(or distinguishability) condition is given by
$\langle v_{i}|e|v_{j}\rangle=0.$ (8)
If $[\bar{e}]=[\mathbf{0}],$ then $e\in C$ due to $d\leq d_{m}.$ Hence, $\exists~{}g\in G$ such that $e=\bar{\gamma}g$ (where $|\gamma|=1$), and we have $\langle v_{i}|e|v_{j}\rangle=\gamma\langle v_{i}|g|v_{j}\rangle=\gamma\lambda_{[j]}(\bar{g})\langle v_{i}|v_{j}\rangle=0.$ If $[\bar{e}]\neq[\mathbf{0}],$ then $0<\|[\bar{e}]\|\leq w_{Q}([\bar{e}])<d.$ Since $[i]$ and $[j]$ are codewords in $\Omega$ with distance $d,$ we have
$[i]\neq[\bar{e}]+[j].$ (9)
Based on the property that eigenvectors with different eigenvalues are orthogonal, it follows that $\langle v_{i}|e|v_{j}\rangle=0.$ So, the necessary condition is satisfied. When $\delta_{ij}=1,$ the measurement condition is
$\langle v_{i}|e|v_{i}\rangle=\langle v_{j}|e|v_{j}\rangle=f(e).$ (10)
If $[\bar{e}]\neq[\mathbf{0}],$ then $f(e)=0.$ If $[\bar{e}]=[\mathbf{0}],$ then $e\in C$ due to $d\leq d_{m}.$ Hence, $\exists~{}g\in G$ such that $e=\bar{\gamma}g.$ To satisfy $\langle v_{i}|e|v_{i}\rangle=\gamma\lambda_{[i]}(\bar{e})=\langle v_{j}|e|v_{j}\rangle=\gamma\lambda_{[j]}(\bar{e})=f(e),$ we need $\lambda_{[i]}(\bar{e})=(-1)^{([i],\bar{e})_{s}}=\lambda_{[j]}(\bar{e})=(-1)^{([j],\bar{e})_{s}}.$ This implies
$[i]-[j]\in([\bar{e}])^{\bot}_{s},$ (11)
and since this relation must hold for every $[\bar{e}]\in C(d-1),$ we have
$[i]-[j]\in C(d-1)^{\bot}_{s}.$ (12)
Therefore, when $\Omega$ belongs to the same coset of $\frac{\mathbb{F}^{2n}_{2}}{C(d-1)^{\bot}_{s}},$ the measurement condition is satisfied. $\square$ Define $Q(\Omega)$ as the Quotient Space Quantum Code ($\mathbf{{QSQC}}$). From the proof, it is clear that formulas (9) and (12) give the necessary and sufficient conditions for choosing invariant subspaces. Circuit Design  First, design $Q([\mathbf{0}])$ using additive codes. Then, obtain $Q(\Omega)$ through the gate circuit of operator $e$ ($[\bar{e}]\in\Omega$). Decoding Method  Suppose the sender sends $|v_{i}\rangle\in Q([i]),$ and the receiver obtains the quantum state $|v_{j}\rangle=e|v_{i}\rangle.$ By applying the generator elements in $G,$ the error syndrome is determined, resulting in $\lambda_{[j]}.$ It is known that $|v_{j}\rangle\in Q([j]).$ Then, by using $\Omega,$ the state $[j]$ is corrected to $[i],$ and $|v_{i}\rangle$ is obtained using the decoding method of additive codes. Unique Advantages  If each state in $Q([i])$ is considered as the same codeword $[i],$ then $Q(\Omega)$ can be viewed as a classical code and it’s immune to $G_{s}^{\bot}.$ Next, I discuss some special cases of Theorem 2. It is easy to see that when $C(d-1)=\{\mathbf{0}\},$ $Q(\Omega)$ becomes a nondegenerate code. When $L=1,$ we obtain additive codes $((n,2^{k},d)).$ When $k=0,$ $C=C_{s}^{\bot}$ and $(C_{s}^{\bot})\setminus C=\emptyset,$ therefore, we can remove the restriction $d\leq d_{m}$ in Theorem 2, resulting in CWS codes $((n,L,d)).$ Since $C(d-1)\subset C\subset C_{s}^{\bot}\subset C(d-1)_{s}^{\bot},$ let $s=\text{dim}(C(d-1))$ and $d_{s}=\min\{w_{Q}(c)\mid c\in(C(d-1)_{s}^{\bot})\setminus C(d-1)\}.$ It is known that the additive code containing $Q(\Omega)$ is $[[n,n-s,d_{s}]].$ Let $B(d)=\{x|w_{H}(x)\leq d\},$ due to $\overline{E}(d-1)\subset B(2d-2),$ if the distance in equation (5) is defined as $\|x\|=w_{H}(x),$ then the distance in Theorem 2 should be adjusted to $12d-$ 1 . New Bounds  New codes have inspired new bounds. The requirement in Theorem 2 that $\Omega$ belongs to $\mathbb{F}^{2n}_{2}/C(d-1)^{\bot}_{s}$ is to satisfy the measurement condition (12). Due to the translational invariance property of distance, it suffices to find equivalent QSC within $C(d-1)^{\bot}_{s}.$ Thus, in contrast to the classical method, in the sphere-packing method the full space to be filled is not $V$ but $C(d-1)^{\bot}_{s},$ the codeword $[i]\in C(d-1)^{\bot}_{s}$ after the action of errors is not $B([i],t)=\{[j]\mid d([j],[i])\leq t\},$ but $MB([i],t)=B([i],t)\bigcap C(d-1)^{\bot}_{s},$ and the quotient sphere centered at the origin is $[E](t)=\{[\bar{e}]\mid\|[\bar{e}]\|\leq t\},$ which can be measured as $ME(t)=[E](t)\bigcap C(d-1)^{\bot}_{s}.$ By the definition of $B([i],t),$ it is easy to see that $B([i],t)=[i]+\ [E](t).$ Let $[\bar{e}]\in[E](t),$ since $C(d-1)^{\bot}_{s}$ is a linear space, we have $[i]+[\bar{e}]\in C(d-1)^{\bot}_{s}\Leftrightarrow[\bar{e}]\in C(d-1)^{\bot}_{s}\Leftrightarrow[\bar{e}]\in ME(t).$ Thus, $|MB([i],t)|=|ME(t)|,$ $|C(d-1)^{\bot}_{s}|=2^{n-k-s},$ where $|A|$ is the number of different cosets in $[A].$Analogous to deriving bounds of classical code, I obtain two types of quantum bounds–measurement bounds.
Theorem 3 (Hamming type)
$Q(\Omega)((n,2^{k}L,d))$ satisfies $2^{k}L|ME(\lfloor\frac{d-1}{2}\rfloor)|\leq 2^{n-s}.$
Theorem 4 (G-V type)
If $2^{k}L|ME(d-1)|<2^{n-s},$ then there exists a $Q(\Omega)((n,2^{k}(L+1),\geq d)).$
When $Q(\Omega)$ is nondegenerate for $d,$ i.e., $C(d-1)^{\bot}_{s}=V,$ the Hamming bound in Theorem 3 reduces to the General Hamming bound [21]. If we can prove that $\sum_{i=0}^{\lfloor\frac{d-1}{2}\rfloor}3^{i}{n\choose i}\leq 2^{s}|ME(\lfloor\frac{d-1}{2}\rfloor)|,$ then the Hamming bound for nondegenerate codes can be applied to degenerate codes. When $s=n-k,$ we can only construct additive codes. Representatives  To proof the quantum Singleton bound [19, 22], we need to choose appropriate representatives for the cosets. For each subspace $H,$ there exists a subspace $U$ such that $V=H\oplus U.$ For any vector $v=h+u\in V,$ I define the projection mapping $\mathcal{P}$ as follows
$\mathcal{P}:V\rightarrow U,\quad\mathcal{P}(v)=\mathcal{P}(h+u)=u.$ (13)
It is clear that $\ker(\mathcal{P})=H,$ and thus $W=V/H\cong U.$ Therefore, each coset $[x]$ in $W$ has a unique representative $\mathcal{P}x,$ which belongs to $U.$ Let $\vec{T}=(h_{1},\dots,h_{n+k},u_{1},\dots,u_{n-k})$ be a basis for $V,$ where $h$ is a basis for $H$ and $u$ is a unit basis for $U.$ The matrix of $\mathcal{P}$ concerning this basis is given by
$P=\mathcal{P}\vec{T}=\begin{pmatrix}\mathbf{0}_{n+k}&\mathbf{0}\\ \mathbf{0}&\mathbf{1}_{n-k}\end{pmatrix}.$ (14)
Let $x=\vec{T}\tilde{x},$ then $Px=(\mathbf{0},\tilde{x}_{n+k+1},\dots,\tilde{x}_{2n})^{T}.$ Thus, the representatives only need to be represented by $n-k$ bits. To extract the last $n-k$ bits of $Px,$ we can still represent $P$ as $\begin{pmatrix}\mathbf{0}&\mathbf{1}_{n-k}\end{pmatrix},$ then $Px=(\tilde{x}_{n+k+1},\dots,\tilde{x}_{2n})^{T}.$ Based on this, I define the quotient projection norm as $\|[x]\|_{p}=w_{H}(Px).$ It’s obvious that $\|[x]\|_{p}$ satisfies the norm axioms, and $\|[x]\|_{p}\geq w_{H}([x])=\inf_{\tilde{x}\in[x]}w_{H}(\tilde{x}).$ If we want to construct QSQC using this norm, adjustments are needed in Theorem 2. I now present the Singleton bound.
Theorem 5 (Singleton)
If $Q(\Omega)[[n,k+l,d]]$ satisfies $n\equiv k\mod 2$ and $l\equiv 0\mod 2,$ then $n\geq k+l+2d-2.$
Proof Since there exists a $\mathbb{F}_{2}$-linear isomorphism $\psi:\mathbb{F}_{2}^{2n}\rightarrow\mathbb{F}_{4}^{n}$ that preserves inner products [7], and $d_{H}(\psi x,\psi y)=d_{Q}(x,y),$ we know that there exists $P=\mathcal{P}\circ\psi:\frac{\mathbb{F}_{2}^{2n}}{C_{s}^{\bot}}\rightarrow\mathbb{F}_{4}^{\frac{n-k}{2}}$ such that $d_{H}(P[x],P[y])\geq d_{Q}([x],[y]).$ Since $\Omega$ has parameters $[n-k,l,d],$ $P\Omega$ can be regarded as a classical code $[\frac{n-k}{2},\frac{l}{2},t]_{4}.$ Let $[i],[j]\in\Omega,$ and let $t=d_{H}(P[i],P[j]).$ Since $d_{H}(P[i],P[j])\geq d_{Q}([i],[j])\geq d,$we have $t\geq d.$ According to the Singleton bound for classical codes, $\frac{n-k}{2}\geq\frac{l}{2}+t-1.$ Hence, $n-k\geq l+2d-2.$square Example  I present four constructions of $((8,8,3))$ codes. First, we select the self-orthogonal code $C_{8},$ also known as the Grassl’s code tables [23]
$C_{8}=\begin{pmatrix}10001011|00101101\\ 01001110|00111010\\ 00101101|01001110\\ 00010111|01011001\\ 00000000|11111111\end{pmatrix}.$
By adding the base vectors
$\begin{pmatrix}\alpha_{1}\\ \alpha_{2}\\ \alpha_{3}\\ \end{pmatrix}=\begin{pmatrix}11101000|00000000\\ 01110100|00000000\\ 11010010|00000000\end{pmatrix},$
we obtain three additional self-orthogonal codes
$C_{81}=\begin{pmatrix}C_{8}\\ \alpha_{1}\end{pmatrix},\quad C_{82}=\begin{pmatrix}C_{81}\\ \alpha_{2}\end{pmatrix},\quad C_{83}=\begin{pmatrix}C_{82}\\ \alpha_{3}\\ \end{pmatrix}.$
All four codes have minimum distance $d_{m}=3$ and are nondegenerate for $d=3.$ Next, we found the following QSCs
$\Omega_{8}(5,1,3)~{}=$$\{[\mathbf{0}]\},$
$\Omega_{81}(6,2,3)=$$\{[\mathbf{0}],[01100000|10010000]\},$
$\Omega_{82}(7,4,3)=$$\{[\mathbf{0}],[10100000|11000000],$
$[01100000|10010000],$ [11000000|01010000]},[11000000|01010000]\},[ 11000000 | 01010000 ] } ,
$\Omega_{83}(8,8,3)=$$\{[\mathbf{0}],[10100000|11000000],$
$[10010000|10100000],$ $[00110000|01100000],$
$[01100000|10010000],$ $[11000000|01010000],$
$[11110000|00110000],$ [01010000|11110000]}.[01010000|11110000]\}.[ 01010000 | 11110000 ] } .
By applying theorem 2, we obtain the following codes: $Q(\Omega_{8})((8,2^{3},3)),$ $Q(\Omega_{81})((8,2^{2}\cdot 2,3)),$ $Q(\Omega_{82})((8,2\cdot 4,3)),$ and $Q(\Omega_{83})((8,8,3)).$ Next, I consider a degenerate CWS code $((12,2,5))$,the self-dual code is given by
$C_{12}=\begin{pmatrix}100001000000|010100011000\\ 010001000000|010111001100\\ 001001000000|010010000110\\ 000101000000|011101100100\\ 000011000000|001100110110\\ 000000100010|001100111010\\ 000000010010|010111110110\\ 000000001010|000101010100\\ 000000000110|010100101000\\ 000000000001|000000000000\\ 000000000000|111111000000\\ 000000000000|000000111110\\ \end{pmatrix},$
with $C(4)=(000000000001|000000000000)$. We found
$\Omega_{12}(15,2,5)=\{[\mathbf{0}],[000000111110|000000000000]\}.$
It is known that $\Omega_{12}\subset C(4)^{\bot}_{s}$,so we obtain the CWS $Q(\Omega_{12})((15,2,5)).$ Discussion  I did not find a simple method to compute the distance of QSC and it is also unclear if there exists a strictly nonadditive and strictly non-CWS code. If we consider each invariant subspace as a stone, additive codes use the largest stone for tiling, while CWS uses the smallest stone for tiling. QSQC encompasses tiling schemes with stones of various sizes. As the volume of the stones decreases, the gaps between them become smaller, but the number of stones to be selected increases. Therefore, intuitively, CWS codes can achieve the optimal parameters but have the highest algorithmic complexity. On the other hand, QSQCs are upgraded to additive codes [24] and can be easily extended to asymmetric, non-binary, self-orthogonal, symplectic self-orthogonal, entanglement-assisted codes, and so on.
Acknowledgements.
I would like to thank Keqin Feng and Hao Chen for the book [14], whose proofs of additive codes inspired my writing of this letter. I was also grateful to Victor V. Albert for the discussion on examples.